Rockbox

Tasklist

FS#6309 - H120 optical out setting doesn't stick

Attached to Project: Rockbox
Opened by Alan Morse Davies (stuffedspacedog) - Thursday, 09 November 2006, 09:31 GMT
Task Type Bugs
Category Configuration
Status Closed
Assigned To No-one
Operating System Iriver H100 series
Severity Low
Priority Normal
Reported Version
Due in Version Undecided
Due Date Undecided
Percent Complete 100%
Votes 0
Private No

Details

With Rockbox in flash on the H120, the setting for optical out is not sticky in the latest builds (8th November), and is set to OFF after each reboot.
This task depends upon

Closed by  Dominik Riebeling (bluebrother)
Thursday, 16 November 2006, 07:39 GMT
Reason for closing:  Fixed
Additional comments about closing:  fixed according to comments. If it\'s still an issue we can reopen this task.
Comment by arpi (arpi) - Sunday, 12 November 2006, 18:17 GMT
I can confirm that and even after a playback stop and playback start the optical out will revert to OFF.
It goes like this:

After switching on the player the optical out LED is ON (without resume)
After playback start the LED if is OFF
I set the optical out to ON, so now the LED is ON again
After playback stop the LED if remains ON
And after playback start the LED if is OFF

By the way:
How can you determine if Rockbox is in flash?

H140, build 20061108
Comment by Davide (Davide-NYC) - Monday, 13 November 2006, 07:23 GMT
Ditto: confirmed.
Comment by Michael Sevakis (MikeS) - Monday, 13 November 2006, 23:52 GMT
I knew about this bug when I comitted the new recording with the intent to fix it quickly. Had to do some target tree moves first before I could do it in the way I wanted. Confirm that is it now fixed (should be, tested it) and I'll close this task.

The optical output should always be on or off per your setting in Playback|Optical Output. You'll still get the dim output if using digital in as the source with the output off. Nothing can be done there.

EDIT: Of course after reading the reported problem (been really scattered around lately :P) the bug that was part of the commit shoudn't have anything at all to do with flashed fw or not. My repair of the digital out behavior was for a known issue of it being left in the wrong state after a recording session but if this fixes the reported problem too all the better.
Comment by Alan Morse Davies (stuffedspacedog) - Tuesday, 14 November 2006, 15:55 GMT
Confirmed, fixed for me.
Comment by Davide (Davide-NYC) - Tuesday, 14 November 2006, 22:32 GMT
I checked this visually only, but seems to be fixed here as well.
I don't access to an optical in this minute, but I presume signal is passing properly.
At the very least the light stays on as it should. :-)
Comment by Michael Sevakis (MikeS) - Tuesday, 14 November 2006, 22:39 GMT
I checked it before committing and the output is correct however I need to mod the cables so I can have two plugged in at once. They're too close together to check a direct digital feedthrough.
Comment by Michael Sevakis (MikeS) - Wednesday, 15 November 2006, 01:13 GMT
Maybe I should consider this bug fixed?
Comment by arpi (arpi) - Thursday, 16 November 2006, 07:11 GMT
Yes, from my part it is now fixed.
Tried, working.
Thank you.

Loading...